OpenStudy (anonymous):

a gardener is planning to make a rectungular garden with an area of 80ft square. she has 12 yards of fencing to put ground the perimeter of the garden. what should the dimension of the garden be?

OpenStudy (anonymous):

is dis question been asked in sum kind of homework? u r the third person in 3 hours asking dis

OpenStudy (anonymous):

yup ... its group work

OpenStudy (anonymous):

can you please help??

OpenStudy (anonymous):

let the to sides be x nd y, so arey=x*y=80 and x+y=12/2=6 yards, now convert it into feets= 18 feets so x=18-y puttting in 1st equation, y(18-y)=80 clearly the asnwer is 8 nd 10

OpenStudy (anonymous):

u understood ?

OpenStudy (anonymous):

yup.. tnx

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

what's ur job??

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

you seem so profesional. =D

OpenStudy (anonymous):

my job is to deal with the best minds of my country, mayb thts y i sound professional